# Multi-LLM-Agent Systems: Techniques and Business Perspectives

### v1

@@ 3. Key Business Aspects
Drawing from our research on MLAS, we present our vision of its business implications across three critical dimensions: privacy preservation, traffic monetization, and intelligence monetization. We anticipate how MLAS could reshape business paradigms and explore the potential trajectories of its commercial applications.
3.1. Privacy Preservation in MLAS
The rise of MLAS introduces new privacy challenges that go beyond those of traditional multi-agent systems. Unlike conventional agents, which exchange structured data, LLM agents handle rich, contextual information that may contain sensitive data embedded in natural language conversations, reasoning processes, and knowledge representations. Privacy preservation in MLAS is critical because these systems process natural language data, which can inadvertently leak sensitive information through semantic connections and implicit knowledge representations.
Privacy-Preserving Challenges. We identify three distinct levels of privacy concerns that require systematic analysis and novel solutions:
@@ 3.1. Privacy Preservation in MLAS
  • At the system architecture level, the distributed nature of MLAS complicates the enforcement of privacy guarantees across all components while maintaining system efficiency. The dynamic interactions between agents and their evolving knowledge further challenge the implementation of robust privacy protections.
@@
At the foundational level, Homomorphic Encryption (HE) (Cheon et al., 2017) enables secure computation on encrypted data, supporting private agent-to-agent communication and inference. While promising in privacy-preserving machine learning (Gilad-Bachrach et al., 2016) and secure data sharing (Chen et al., 2017), HE’s computational complexity remains a significant challenge in MLAS.
Secure Multi-Party Computation (SMPC) (Lindell, 2020) enables secure collaborative computation among multiple agents. SMPC has been used in privacy-preserving data analysis (Li and Xu, 2019) and collaborative learning (Knott et al., 2021) in traditional multi-agent systems, but scaling this technology to large MLAS remains known.
For system-level protection, Trusted Execution Environments (TEEs) (Costan, 2016) provide hardware-based security guarantees. Technologies such as Intel SGX (Costan, 2016), ARM TrustZone (Pinto and Santos, 2019), and AMD SEV (Sev-Snp, 2020) create secure enclaves for sensitive computations. However, integrating TEEs into MLAS requires careful consideration of security and performance trade-offs.
@@
Research Directions and Open Challenges. Advancing privacy preservation in MLAS involves addressing several key issues. First, existing privacy metrics do not fully capture the complexities of semantic information leakage in natural language processing. There is a need for MLAS-specific privacy frameworks that account for both direct and indirect information flows in semantic spaces. Second, integrating privacy-preserving technologies across MLAS requires a unified approach that combines data protection, secure computation, and communication security. Performance optimization and scalability remain major hurdles, especially as agent networks expand, and maintaining privacy while enabling efficient collaboration is critical. Future research should focus on creating comprehensive privacy frameworks tailored to MLAS. This includes standardizing privacy protocols, developing efficient implementations, and establishing evaluation metrics to ensure the practical deployment of privacy-preserving MLAS systems.
@@ 3.2. Traffic monetization
Traffic Monetization in MLAS involves generating commercial value by managing user traffic and optimizing ads, using the strengths of various agents. This includes improving traffic flow, boosting click-through rates (CTR), and increasing conversion rates (CVR) (Kumar and Reinartz, 2016). MLAS leverages each agent’s capabilities to enhance user engagement and make advertising strategies more effective. The system also ensures fair and transparent revenue allocation. This section explores Traffic Monetization in MLAS, covering business scenarios, revenue models, profit allocation, and the roles of application agents.
@@
Business Scenarios and Revenue Generation. In MLAS, agents analyze user behaviors and preferences to optimize traffic management and deploy targeted advertisements. Through real-time data analytics, agents identify the most relevant ads to engage users, thereby enhancing click-through rates (CTR) and conversion rates (CVR). This approach builds user profiles and uses intelligent recommendation systems to personalize ads, boosting engagement and driving purchases. Revenue in MLAS mainly comes from advertising, using Cost Per Click (CPC) and Cost Per Action (CPA) models (Kannan and Li, 2017). In the CPC model, advertisers pay based on clicks, with agents earning commissions based on their contribution to traffic management and ad effectiveness. In the CPA model, payments are made for completed purchases, rewarding agents who drive conversions with a higher share of the revenue. In addition to advertising, income can be generated through user subscriptions for premium features or personalized services within specific applications, such as advanced analytics dashboards or exclusive access to specialized tools. These monetization strategies work together, with agents optimizing traffic flow, user engagement, and ad targeting to drive overall profitability.
Profit Allocation Mechanisms. Converting revenue into profits that are fairly distributed among agents is key to Traffic Monetization. The process starts by assessing each agent’s contribution to traffic generation, ad clicks, and conversions, using metrics like CTR and CVR to quantify individual impact. To ensure fairness, MLAS may use blockchain-based smart contracts to automate the distribution process, minimizing bias and human error. Additionally, a scoring system rates agents based on their performance, including factors like user feedback and engagement. Agents with higher scores receive a larger share of revenue, incentivizing better performance and continuous optimization. Attribution methods, such as the Shapley Value, ensure profits are allocated based on each agent’s contribution to the system (Shapley, 1953). Dynamic adjustment mechanisms allow real-time updates to revenue shares based on agent performance and market conditions, ensuring fair compensation for optimizing traffic and ads. Incorporating metrics like CPM (Cost Per Mille) adds a more nuanced approach to revenue allocation, offering a deeper understanding of ad performance beyond just clicks and conversions (Brynjolfsson and McAfee, 2014).
@@
Traffic Monetization in MLAS creates a system where revenue is generated and shared through the collaboration of different agents. By improving traffic management, optimizing ads, and ensuring fair distribution, MLAS makes sure all agents benefit fairly. Future research should focus on improving attribution models like the Shapley Value to make profit allocation even fairer. Adding metrics like CPM will also help allocate revenue more accurately and improve monetization strategies.
@@ 3.3. Intelligence Monetization
Intelligence Monetization in Multi-LLM-Agent Systems represents a significant evolution in AI commercialization by leveraging the collaborative capabilities of specialized agents. Unlike traditional single-model paradigms, multi-agent systems enable dynamic interactions among specialized agents, each designed to address specific tasks, thereby facilitating the creation of more versatile and robust intelligence solutions. This paradigm has shown promise in various applications, as exemplified by Microsoft’s Copilot Studio Platform, launched on November 19, 2024. The platform supports an ecosystem of over 1,800 large models and offers open APIs and integration tools, enabling enterprises to incorporate agent technology into workflows and applications for enhanced customization and scalability (VentureBeat, 2024).
@@
Revenue Generation through Data-Driven Services. A key revenue model in Intelligence Monetization within MLAS is the sale of data-driven services (Sjödin et al., 2021; Faroukhi et al., 2020; Hajipour et al., 2023; VentureBeat, 2024). Specialized agents analyze distinct datasets—such as consumer preferences, product usage, and market trends—to generate actionable insights. These insights are delivered in the form of reports, forecasts, or tailored recommendations that businesses can purchase. For instance, one agent may provide personalized user behavior insights to enhance marketing, while another offers market trend analysis to guide strategic planning. By integrating specialized agents, MLAS offers a broad range of insights, which can be monetized through subscription models or one-time reports. In practice, successful implementations like OpenAI’s GPT-4 API (OpenAI et al., 2024) have demonstrated how multiple specialized models can work in concert, with distinct agents handling different aspects of the intelligence pipeline. These include specialized agents for data processing and preprocessing, deep pattern recognition and insight generation, recommendation transformation, and platform integration.
@@
Innovative Licensing and Agent Marketplaces. MLAS also introduces novel licensing approaches that move beyond traditional software licensing. One of the most prominent is the Agent-as-a-Service (AaaS) (Rajaei, 2024), as seen in Google Cloud’s AutoML, which enables dynamic agent deployment based on computational needs, with usage-based pricing and automatic scaling (Cloud, 2024). Complementing this will be the emergence of agent marketplace platforms, creating ecosystems for third-party agent development and deployment, as demonstrated by Hugging Face’s model hub adapted for deployment of LLMs (Face, 2024). Furthermore, hybrid deployment architectures will become increasingly popular, combining on-premise agent deployment for sensitive operations with cloud-based agents for scalable tasks, as seen in IBM’s Watson services which use distributed agent architecture (IBM, 2024).
@@
Intelligence Monetization within MLAS offers a sustainable revenue framework by harnessing the collective intelligence of specialized agents. Looking ahead, as demand for AI-driven insights grows, MLAS’s role in delivering scalable, actionable intelligence will continue to expand. Future developments should focus on enhancing agent collaboration for real-time insights and adapting to emerging business models and industries, offering tailored solutions to meet specific needs.
@@ 3.4. Integration of 3 Business Aspects
The three key business aspects of MLAS form an interconnected framework that drives both commercial success and ethical operation. Data privacy ensures trust and compliance while allowing secure data usage. This foundation supports traffic monetization by generating user engagement data while adhering to privacy regulations. Intelligence monetization transforms these privacy-preserved interactions into actionable insights and services. As MLAS evolves, maintaining a balance between privacy, commercial success, and technological progress will be critical for long-term sustainability. Future development should focus on strengthening these connections while adapting to evolving privacy regulations and market demands.
Figure 4. Architectures of MLAS.
@@ 4. Case Study
Building on the technical foundations and business considerations of MLAS, we now delve into real-world implementations to illustrate how these theoretical frameworks are realized in practice. Through carefully selected case studies, we explore how different architectural choices influence system efficiency, data privacy, and monetization capabilities. These examples not only validate our theoretical insights but also shed light on the challenges and opportunities inherent in deploying MLAS solutions.
4.1. Architectures in MLAS
The implementation of MLAS in real-world applications reveals various architectural patterns, each designed to address specific operational requirements and constraints.
Figure 5. Centralized Architecture of MLAS.
Figure 6. Centralized data handling of MLAS.
As we analyze current MLAS deployments, we observe that architectural choices significantly influence system capabilities, from privacy protection to operational efficiency. Figure 4 illustrates four fundamental patterns that have emerged in practice:
@@ 4.1. Architectures in MLAS
4.2. A Decentralized Star Architecture in MLAS
Figure 5 and Figure 6 present our first case study of MLAS implementation in a music service scenario. The system uses a centralized architecture where several agents, including a Personal Agent, an Orchestrator Agent, and a Song Agent, collaborate to process user’s music playback requests. In this setup, the Orchestrator Agent acts as a central hub, managing communication and coordinating tasks among the agents.
@@ 4.2. A Decentralized Star Architecture in MLAS
Figure 7. Decentralized Star Architecture of MLAS.
Figure 8. Decentralized data handling of MLAS.
@@ 5. Conclusion & Future
In this paper, we provide our analysis on the future development of multi-LLM-agent systems (MLAS) from the perspectives of techniques and business. Technically, compared to the traditional single-LLM-agent systems, MLAS has a higher potential for overall performance and system flexibility; and commercially, MLAS brings the feasibility of proprietary data preservability and monetization through traffic and intelligence, which essentially incentivizes various entities to contribute to the whole ecosystem. Several effective protocols for multi-agent communication and collaboration are developed in progress, which will drive the implementation of the MLAS ecosystem towards achieving artificial collective intelligence in the near future.

### v2

@@ 3. Key Business Aspects
Drawing from our research on LaMAS, we present our vision of its business implications across three critical dimensions: privacy preservation, traffic monetization, and intelligence monetization. We anticipate how LaMAS could reshape business paradigms and explore the potential trajectories of its commercial applications.
3.1. Privacy Preservation in LaMAS
The rise of LaMAS introduces new privacy challenges that go beyond those of traditional multi-agent systems. Unlike conventional agents, which exchange structured data, LLM agents handle rich, contextual information that may contain sensitive data embedded in natural language conversations, reasoning processes, and knowledge representations. Privacy preservation in LaMAS is critical because these systems process natural language data, which can inadvertently leak sensitive information through semantic connections and implicit knowledge representations.
Privacy-Preserving Challenges. We identify three levels of privacy concerns that require systematic analysis and novel solutions:
@@ 3.1. Privacy Preservation in LaMAS
  • At the system architecture level, the distributed nature of LaMAS complicates the enforcement of privacy guarantees across all components while maintaining system efficiency. The dynamic interactions between agents and their evolving knowledge further challenge the implementation of robust privacy protections.
@@
At the foundational level, Homomorphic Encryption (HE) (Cheon et al., 2017) enables secure computation on encrypted data, supporting private agent-to-agent communication and inference. While promising in privacy-preserving machine learning (Gilad-Bachrach et al., 2016) and secure data sharing (Chen et al., 2017), HE’s computational complexity remains a significant challenge in LaMAS.
Secure Multi-Party Computation (SMPC) (Lindell, 2020) enables secure collaborative computation among multiple agents. SMPC has been used in privacy-preserving data analysis (Li and Xu, 2019) and collaborative learning (Knott et al., 2021) in traditional multi-agent systems, but scaling this technology to large LaMAS remains known.
For system-level protection, Trusted Execution Environments (TEEs) (Costan, 2016) provide hardware-based security guarantees. Technologies such as Intel SGX (Costan, 2016), ARM TrustZone (Pinto and Santos, 2019), and AMD SEV (Sev-Snp, 2020) create secure enclaves for sensitive computations. However, integrating TEEs into LaMAS requires careful consideration of security and performance trade-offs.
@@
Research Directions and Open Challenges. Advancing privacy preservation in LaMAS involves addressing several key issues. First, existing privacy metrics do not fully capture the complexities of semantic information leakage in natural language processing. There is a need for LaMAS-specific privacy frameworks that account for both direct and indirect information flows in semantic spaces. Second, integrating privacy-preserving technologies across LaMAS requires a unified approach that combines data protection, secure computation, and communication security. Performance optimization and scalability remain major hurdles, especially as agent networks expand, and maintaining privacy while enabling efficient collaboration is critical. Future research should focus on creating comprehensive privacy frameworks tailored to LaMAS. This includes standardizing privacy protocols, developing efficient implementations, and establishing evaluation metrics to ensure the practical deployment of privacy-preserving LaMAS systems.
@@ 3.2. Traffic Monetization
Traffic Monetization in LaMAS involves generating commercial value by managing user traffic and optimizing ads, using the strengths of various agents. This includes improving traffic flow, boosting click-through rates (CTR), and increasing conversion rates (CVR) (Kumar and Reinartz, 2016). LaMAS leverages each agent’s capabilities to enhance user engagement and make advertising strategies more effective. The system also ensures fair and transparent revenue allocation. This section explores Traffic Monetization in LaMAS, covering business scenarios, revenue models, profit allocation, and the roles of application agents.
@@
Business Scenarios and Revenue Generation. In LaMAS, agents analyze user behaviors and preferences to optimize traffic management and deploy targeted advertisements. Through real-time data analytics, agents identify the most relevant ads to engage users, thereby enhancing click-through rates (CTR) and conversion rates (CVR). This approach builds user profiles and uses intelligent recommendation systems to personalize ads, boosting engagement and driving purchases. Revenue in LaMAS mainly comes from advertising, using Cost Per Click (CPC) and Cost Per Action (CPA) models (Kannan and Li, 2017). In the CPC model, advertisers pay based on clicks, with agents earning commissions based on their contribution to traffic management and ad effectiveness. In the CPA model, payments are made for completed purchases, rewarding agents who drive conversions with a higher share of the revenue. In addition to advertising, income can be generated through user subscriptions for premium features or personalized services within specific applications, such as advanced analytics dashboards or exclusive access to specialized tools. These monetization strategies work together, with agents optimizing traffic flow, user engagement, and ad targeting to drive overall profitability.
Profit Allocation Mechanisms. Converting revenue into profits that are fairly distributed among agents is key to Traffic Monetization. The process starts by assessing each agent’s contribution to traffic generation, ad clicks, and conversions, using metrics like CTR and CVR to quantify individual impact. To ensure fairness, LaMAS may use blockchain-based smart contracts to automate the distribution process, minimizing bias and human error. Additionally, a scoring system rates agents based on their performance, including factors like user feedback and engagement. Agents with higher scores receive a larger share of revenue, incentivizing better performance and continuous optimization. Attribution methods, such as the Shapley Value, ensure profits are allocated based on each agent’s contribution to the system (Shapley, 1953). Dynamic adjustment mechanisms allow real-time updates to revenue shares based on agent performance and market conditions, ensuring fair compensation for optimizing traffic and ads. Incorporating metrics like CPM (Cost Per Mille) adds a more nuanced approach to revenue allocation, offering a deeper understanding of ad performance beyond just clicks and conversions (Brynjolfsson and McAfee, 2014).
@@
Traffic Monetization in LaMAS creates a system where revenue is generated and shared through the collaboration of different agents. By improving traffic management, optimizing ads, and ensuring fair distribution, LaMAS makes sure all agents benefit fairly. Future research should focus on improving attribution models like the Shapley Value to make profit allocation even fairer. Adding metrics like CPM will also help allocate revenue more accurately and improve monetization strategies.
@@ 3.3. Intelligence Monetization
Intelligence Monetization in LLM-based Multi-Agent Systems represents a significant evolution in AI commercialization by leveraging the collaborative capabilities of specialized agents. Unlike traditional single-model paradigms, multi-agent systems enable dynamic interactions among specialized agents, each designed to address specific tasks, thereby facilitating the creation of more versatile and robust intelligence solutions. This paradigm has shown promise in various applications, as exemplified by Microsoft’s Copilot Studio Platform, launched on November 19, 2024. The platform supports an ecosystem of over 1,800 large models and offers open APIs and integration tools, enabling enterprises to incorporate agent technology into workflows and applications for enhanced customization and scalability (VentureBeat, 2024).
@@
Revenue Generation through Data-Driven Services. A key revenue model in Intelligence Monetization within LaMAS is the sale of data-driven services (Sjödin et al., 2021; Faroukhi et al., 2020; Hajipour et al., 2023; VentureBeat, 2024). Specialized agents analyze distinct datasets—such as consumer preferences, product usage, and market trends—to generate actionable insights. These insights are delivered in the form of reports, forecasts, or tailored recommendations that businesses can purchase. For instance, one agent may provide personalized user behavior insights to enhance marketing, while another offers market trend analysis to guide strategic planning. By integrating specialized agents, LaMAS offers a broad range of insights, which can be monetized through subscription models or one-time reports. In practice, successful implementations like OpenAI’s GPT-4 API (OpenAI et al., 2024) have demonstrated how multiple specialized models can work in concert, with distinct agents handling different aspects of the intelligence pipeline. These include specialized agents for data processing and preprocessing, deep pattern recognition and insight generation, recommendation transformation, and platform integration.
@@
Innovative Licensing and Agent Marketplaces. LaMAS also introduces novel licensing approaches that move beyond traditional software licensing. One of the most prominent is the Agent-as-a-Service (AaaS) (Rajaei, 2024), as seen in Google Cloud’s AutoML, which enables dynamic agent deployment based on computational needs, with usage-based pricing and automatic scaling (Cloud, 2024). Complementing this will be the emergence of agent marketplace platforms, creating ecosystems for third-party agent development and deployment, as demonstrated by Hugging Face’s model hub adapted for deployment of LLMs (Face, 2024). Furthermore, hybrid deployment architectures will become increasingly popular, combining on-premise agent deployment for sensitive operations with cloud-based agents for scalable tasks, as seen in IBM’s Watson services which use distributed agent architecture (IBM, 2024).
@@
Intelligence Monetization within LaMAS offers a sustainable revenue framework by harnessing the collective intelligence of specialized agents. Looking ahead, as demand for AI-driven insights grows, LaMAS’s role in delivering scalable, actionable intelligence will continue to expand. Future developments should focus on enhancing agent collaboration for real-time insights and adapting to emerging business models and industries, offering tailored solutions to meet specific needs.
@@ 3.4. Integration of 3 Business Aspects
The three key business aspects of LaMAS form an interconnected framework that drives both commercial success and ethical operation. Data privacy ensures trust and compliance while allowing secure data usage. This foundation supports traffic monetization by generating user engagement data while adhering to privacy regulations. Intelligence monetization transforms these privacy-preserved interactions into actionable insights and services. As LaMAS evolves, maintaining a balance between privacy, commercial success, and technological progress will be critical for long-term sustainability. Future development should focus on strengthening these connections while adapting to evolving privacy regulations and market demands.
Figure 4. Architectures of LaMAS.
@@ 4. Case Study
Building on the technical foundations and business considerations of LaMAS, we now delve into real-world implementations to illustrate how these theoretical frameworks are realized in practice. Through carefully selected case studies, we explore how different architectural choices influence system efficiency, data privacy, and monetization capabilities. These examples not only validate our theoretical insights but also shed light on the challenges and opportunities inherent in deploying LaMAS solutions.
4.1. Architectures in LaMAS
The implementation of LaMAS in real-world applications reveals various architectural patterns, each designed to address specific operational requirements and constraints.
Figure 5. Centralized Architecture of LaMAS.
Figure 6. Centralized data handling of LaMAS.
As we analyze current LaMAS deployments, we observe that architectural choices significantly influence system capabilities, from privacy protection to operational efficiency. Figure 4 illustrates four fundamental patterns that have emerged in practice:
@@ 4.1. Architectures in LaMAS
4.2. A Decentralized Star Architecture in LaMAS
Figure 5 and Figure 6 present our first case study of LaMAS implementation in a music service scenario. The system uses a centralized architecture where several agents, including a Personal Agent, an Orchestrator Agent, and a Song Agent, collaborate to process user’s music playback requests. In this setup, the Orchestrator Agent acts as a central hub, managing communication and coordinating tasks among the agents.
@@ 4.2. A Decentralized Star Architecture in LaMAS
Figure 7. Decentralized Star Architecture of LaMAS.
Figure 8. Decentralized data handling of LaMAS.
@@ 5. Conclusion & Future
In this paper, we provide our analysis on the future development of LLM-based Multi-Agent Systems (LaMAS) from the perspectives of techniques and business. Technically, compared to the traditional single-LLM-agent systems, LaMAS has a higher potential for overall performance and system flexibility; and commercially, LaMAS brings the feasibility of proprietary data preservability and monetization through traffic and intelligence, which essentially incentivizes various entities to contribute to the whole ecosystem. Several effective protocols for multi-agent communication and collaboration are developed in progress, which will drive the implementation of the LaMAS ecosystem towards achieving artificial collective intelligence in the near future.